OpenStudy (anonymous):

How would I do this? [-2(12-3)/3]^2

OpenStudy (anonymous):

Parenthesis - (12-3)= 9 Exponent - 2 9^2?

OpenStudy (anonymous):

you start from the innermost parenthesis (which in this case is 12-3). The follow MDAS next (multiply then divide). Then multiply it to itself. =)

OpenStudy (anonymous):

first solve the round bracket and then solve the square bracket (the outer bracket)

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

first part is right. after that you multiply it by -2 then divide it by 3. that is when you raise it to 2.

OpenStudy (anonymous):

its easy. the round bracket is the inner bracket, u have to solve first at all times

OpenStudy (anonymous):

\[\left( \frac{ -2(9) }{ 3 } \right)^{2}\]

OpenStudy (anonymous):

so, (12-3)/3 = 3. did u get that

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

huh..

OpenStudy (anonymous):

then u r left with -2 (3) with square root of 2

OpenStudy (anonymous):

solve -2(3)=-6^2 = 36

OpenStudy (anonymous):

umm... hello